OpenStudy (lilsis76):

Exponential problem: how do I solve this problem? xe^x -2e^x = 0

OpenStudy (lilsis76):

\[xe^x-2e^x=0\]

OpenStudy (anonymous):

\(e^x\) is never zero so factor it out, then ignore it

OpenStudy (anonymous):

Factor out e^x, e^x(x - 2) = 0 Set each factor equal to 0, you get e^x = 0 which has no solution, e^x can never equal 0. And the other factor x - 2 = 0, you get x = 2. Final answer, x = 2.

OpenStudy (lilsis76):

okay YAY! okay I did it right, i was also wondering about the steps if I did it correctly. THANK YOU SO MUCH @Easyaspi314 !!!